# Defense Against Prompt Injection Attack by Leveraging Attack Techniques

### v2

@@ 1 Introduction
In this paper, we propose prompt injection defense methods based on several effective prompt engineering attack techniques. To explain our motivation, consider the example in Figure 1 (a). In this scenario, the malicious prompt (highlighted in red) embedded in the retrieved results consists of an attack prompt followed by an injected instruction. The attack prompt misleads the LLM into ignoring the original input instruction, whose answer could otherwise raise the user’s suspicion. The response to the injected instruction fulfills the attacker’s malicious intent. In contrast, our defense goal is for the LLM to ignore the injected instruction and instead respond to the original input instruction. Interestingly, the defense and attack share similar design goals: inducing the LLM to ignore the unwanted instructions and instead to execute the wanted instructions. This raises an intuitive question: Could attack techniques be repurposed or adapted to develop more robust defense methods? Figure 1 (b) demonstrates how we develop our defense strategy based on the attack techniques: we preserve the attack prompt as the shield prompt, and replace the injected instruction with the original input instruction. We apply this approach with several attack techniques. Moreover, we additionally find that when attackers get access to the conversation template, they can pretend to be the assistant to answer the original input instructions, and then act as the user to request the LLM to answer their injected instruction, posing a serious threat. Inspired by this, we design our defense by acting as the assistant who detects the attack and then acting as the user to confirm the instruction.
We conduct comprehensive experiments to evaluate the effectiveness of our defense methods against various prompt injection attack methods. The results demonstrate that our methods outperform existing training-free defense approaches and are even comparable to fine-tuning-based methods, providing more robust protection. Notably, the defense method based on the most effective attack technique performs the best, reducing the attack success rate (ASR) to nearly zero in certain scenarios. Our contributions are summarized as follows:
@@ 4.2.4 Fake Completion Defense with Template
Table 1: The results of our defense methods compared to baselines against various attack methods in the direct prompt injection scenario. Bold indicates the best performance. All results are reported in %.
@@ 5.1 Experimental Settings
We select popular and strong open-source LLMs as victim models for our experiments. Specifically, we choose Llama3.1-8b-Instruct Dubey et al. (2024), Qwen2-7b-Instruct Yang et al. (2024), and Llama3-8b-Instruct AI@Meta (2024). Throughout the experiments, unless otherwise specified, “Llama3,” “Llama3.1,” and “Qwen2” refer to Llama3-8b-Instruct, Llama3.1-8b-Instruct, and Qwen2-7b-Instruct, respectively.
Evaluation Metrics.
For the security metric, we follow the evaluation protocol of Chen et al. (2024), using the attack success rate (ASR) to assess the effectiveness of the defense methods. We detecte if the answer to the injected instruction appears in the generated response. For the utility metric, we use accuracy to evaluate the potential negative impact of defense methods on model performance. Specifically, we employ the filtered QA dataset Li et al. (2023), which is not attacked and includes the defense mechanism. We request the LLMs to answer the questions and verify whether the correct (golden) answers appear in the responses.
@@
Table 2: The results of our defense methods compared to baselines against various attack methods in the indirect prompt injection scenario. Bold indicates the best performance. All results are reported in %.
@@ 5.3.1 Defense against Direct Attack
We perform the direct prompt injection attack following the approach of Chen et al. (2024), using 208 samples from AlpacaFarm. Table 1 presents the effectiveness of our defense methods in the direct prompt injection scenario. The results show that our methods, which are based on attack techniques, outperform the baselines, regardless of the attack method or the victim model. Among the baseline methods, the Sandwich method performs better on average than the others. The key difference between Sandwich and the other baselines lies in the position of the defense prompt: Sandwich places the defense prompt at the end of the data, similar to our methods. This suggests that placing the defense prompt at the end may interfere with the attack and enhance the defense’s effectiveness. When comparing the victim models, we find that Qwen2 is more vulnerable to attacks, possibly indicating that Qwen2 is less capable of distinguishing between different instructions.
@@ 5.3.2 Defense against Indirect Attack
In addition to evaluating defense against direct prompt injection attacks, we also assess its effectiveness against indirect attacks. The key difference between direct and indirect prompt injection attacks is that, in the case of indirect attacks, the input data is retrieved from external tools, such as search engines, and users are often unaware of the attack. Typically, the retrieved data is longer than direct user input. To evaluate indirect prompt injection attacks, we use the filtered QA dataset from Li et al. (2023). Table 2 shows the results of our defense methods compared to the baselines in the indirect scenario. Our methods continue to outperform the baselines by a significant margin. When comparing both direct and indirect prompt injection attacks, it appears that indirect attacks are easier to defend against. Furthermore, Qwen2 remains the most susceptible model to attacks compared to the other two models.
@@ 5.3.3 Model Utility
A key evaluation metric for defense methods is their potential impact on the model’s utility. To assess the impact of our method, we use the filtered QA dataset from Li et al. (2023). For simplicity, we do not introduce any attacks into the retrieved data content, and we only verify whether the correct (golden) answer appears in the model’s response, with different defense methods. Table 3 presents the utility performance of various defense strategies. Interestingly, most defense strategies do not significantly affect the model’s utility. Moreover, our proposed defense methods can even improve the performance in some scenarios.
@@
Table 3: The model utility when applied with different defense methods. All the results are reported in %.
5.3.4 Model Efficiency
Since LLMs are ultimately integrated into various applications, it is crucial to assess the overhead introduced by the defense method, particularly in terms of generating longer, but less informative, responses. To evaluate efficiency, we use the same QA dataset employed for assessing model utility. As demonstrated in Section 5.3.3, the utility results remain almost consistent across different defense strategies. Therefore, any additional overhead indicates an increase in uninformative content within the responses. Table 8 presents the efficiency of various defense methods. The average overhead of our defense method based on "Fakecom-t" and the defense based on escape attack is slightly higher compared to the baseline with no defense, but it remains relatively low. The overhead for the other defense methods is nearly the same.
@@ 5.4 Ablation Study
In this section, we address several questions regarding our defense methods and the evaluation metrics. We perform comprehensive experiments to solidify the validity and robustness of our approach.
@@
To further validate the effectiveness of our methods, we apply our methods to the closed-source models "GPT-3.5-Turbo" Jiang et al. (2023) and "GPT-4o-Latest" Hurst et al. (2024). Because we cannot change the conversation template, we only compare our methods based on "Ignore attack" and "Fakecom attack" with the baselines against direct prompt injection attack. Table 4 shows the results. From the table we can find out that our methods are also effective on closed-source models, surpassing the previous defense baselines. What’s more, comparing the defense performance of the two models with our defense methods reveals that stronger model is more suitable to our methods, making our methods more applicable.
Defense Methods GPT-3.5-Turbo GPT-4o-Latest
None 32.69 50.48 32.69 88.46 87.50 65.86 92.78 63.46 100.00 100.00
Sandwich 13.94 17.30 8.65 4.32 42.30 15.86 29.32 10.09 5.76 37.98
Reminder 11.05 10.57 10.96 9.61 26.92 14.90 27.88 17.30 89.42 82.69
Isolation 22.59 39.42 24.51 43.26 77.40 52.40 83.17 52.88 94.71 99.51
Spotlight 16.34 31.73 13.46 15.38 71.15 19.71 45.67 15.38 47.11 68.75
Ours-Ignore 2.88 3.36 1.44 0.48 4.32 0.90 0.90 0.40 0.0 0.0
Ours-Fakecom 5.57 12.01 1.44 0.0 14.90 7.21 2.88 5.76 0.90 7.21
Table 4: The results of our defense methods compared with defense baselines applied on closed-source models. Bold indicates the best performance. All results are reported in %.
@@
Beyond prompt-engineering-based attacks, we also examine the effectiveness of our defense methods against the gradient-based attack. We implement direct prompt injection attack with GCG attack method using Llama3, following the implementation by Chen et al. (2024), and construct the suffix with knowledge of the conversation template. Since the GCG suffix is transferable in jailbreak attacks Zou et al. (2023), we also test its transferability in prompt injection attacks using Llama3.1. Table 6 presents the defense results. The first notable observation is that the suffix transferability is limited. Without any defense, the ASR remains relatively low, similar to a naive attack. Comparing with the baselines, our defense methods more effectively mitigate the attack. The method based on “Fakecom-t” is the most effective, reducing the ASR to below 10%.
@@
Although it’s very unlikely for the attacker to be aware of the conversation template—since application providers typically filter out template tokens—we are still interested in assessing the potential harm of such an attack. We utilize the AlpacaFarm for evaluation and table 5 presents the results. The table shows that the fake completion attack with a conversation template can be harmful, and most baseline methods are ineffective. Our methods, which rely on ignoring the attack and using the fake completion strategy, function as intended but result in only a limited decrease in ASR. Our method based on this attack ("Fakecom-t") is effective, and this phenomenon raises our question: Would the effectiveness of the attack methods determine the defense methods designed on them?
@@
Table 5: The results demonstrate how harmful the fake completion attack with the conversation template is. All the results are reported in %.
@@
In previous introduction, we argue that fine-tuning-based methods require significant computational resources. But the fine-tuning-based methods are more effective than previous prompt-engineering-based methods. Therefore, we compare our methods with StruQ Chen et al. (2024), which incorporates prompt injection attack methods into the clean data for fine-tuning. We incorporate the "Naive attack" and the "Ignore attack" respectively for evaluation. Table 7 shows the results. It’s obvious that the ability of StruQ to generalize to the unknown attacks is not satisfactory. Because "Ignore attack" is a part of "Combined attack", "StruQ-Ignore" can defend against "Combined attack" successfully. The generalization ability of our methods is much better, effectively defending against different attacks.
Defense Methods Llama3 Llama3.1
None 87.01 47.11
Sandwich 19.23 22.59
Instructional 28.84 37.50
Reminder 24.51 32.21
Isolation 40.38 39.42
Spotlight 19.71 29.80
Ours-Ignore 12.01 11.05
Ours-Escape 19.23 21.15
Ours-Fakecom 13.94 22.59
Ours-Fakecom-t 9.61 6.73
Table 6: The performance of the defense methods against the gradient-based attack. Bold indicates the best performance. All results are reported in %.
@@ 5.5 Case Study
Figure 16 provides two examples of responses without defense against the ignore and fake completion attacks, both generated by Llama3. Additionally, we include the response generated by the model with defense method based on "Fakecom-t", as it is the most effective defense approach. From the examples, we observe that the ignore attack does not consistently persuade the LLM to ignore the original input instruction and the model may end up executing both instructions. While in this instance, the fake completion attack successfully misleads the LLM to execute the injected instruction directly, this strategy does not always work—there are cases where the model executes both instructions, explaining the failures of the defense methods based on these attacks. In the case of the defense method based on "Fakecom-t", we can observe that the defense method successfully enables the LLM to bypass the injected instruction. What’s more, the response still remains relevant to the original task, suggesting the defense method has little damage on the utility of the model.
@@ 6 Conclusion
7 Limitations
In this paper, we propose defense methods derived from existing attack strategies. However, our focus has been limited to prompt-engineering-based attacks, and we have not explored to design defenses with gradient-based attacks. Generally, gradient-based attacks tend to be more effective. Based on our findings that strong attack methods often inspire strong defense mechanisms, we believe that defenses designed with gradient-based attack methods would likely be more robust. Investigating how to develop such defenses will be a key direction for our future work.
8 Ethical consideration
@@ Appendix A Appendix
A.1 Defense Baselines
@@
A.2 Implementation Details.
A.3 Dataset Additional Information
@@
None 10.55 53.35 88.25 75.30 86.00
StruQ-Naive 0.50 0.60 2.20 35.55 27.30
StruQ-Ignore 0.05 0.05 8.00 35.70 0.05
Ours-Ignore 0.05 0.35 0.30 0.10 1.35
Table 7: Defense performance of our methods and the fine-tuning method StruQ. "StruQ-Naive" means StruQ incorporates the "Naive attack" for fine-tuning. Bold indicates the best performance. All results are reported in %.
@@
None 0.645 0.699 1.184
Sandwich 0.632 0.701 1.173
Instructional 0.630 0.701 1.169
Reminder 0.632 0.701 1.179
Isolation 0.631 0.702 1.171
Spotlight 0.644 0.734 1.193
Ours-Ignore 0.632 0.701 1.171
Ours-Escape 0.644 0.729 1.548
Ours-Fakecom 0.631 0.701 1.171
Ours-Fakecom-t 0.766 0.704 1.200
Table 8: The time cost across various defense methods. All the results are reported in seconds/item.

### v3

@@ 1 Introduction
In this paper, we propose prompt injection defense methods based on several effective prompt engineering attack techniques. To explain our motivation, consider the example in Figure 1 (a). In this example, the malicious prompt (highlighted in red) embedded in the retrieved results consists of an attack prompt followed by an injected instruction. The attack prompt misleads the LLM into ignoring the original input instruction, whose answer could otherwise raise the user’s suspicion. The response to the injected instruction fulfills the attacker’s malicious intent. In contrast, our defense goal is for the LLM to ignore the injected instruction and instead respond to the original input instruction. Interestingly, the defense and attack share similar design goals: inducing the LLM to ignore the unwanted instructions and instead to execute the wanted instructions. This raises an intuitive question: Could attack techniques be repurposed or adapted to develop more robust defense methods? Figure 1 (b) demonstrates how we develop our defense strategy based on the attack techniques: we preserve the attack prompt as the shield prompt, and replace the injected instruction with the original input instruction. We apply this approach with several attack techniques. Moreover, we additionally find that when attackers get access to the conversation template, they can pretend to be the assistant to answer the original input instructions, and then act as the user to request the LLM to answer their injected instruction, posing a serious threat. Inspired by this, we design our defense by acting as the assistant who detects the attack and then acting as the user to confirm the instruction.
We conduct comprehensive experiments to evaluate the effectiveness of our defense methods against various prompt injection attack methods. The results demonstrate that our methods outperform existing training-free defense approaches against both prompt-engineering-based and gradient-based attack methods. Moreover, our methods are even comparable to fine-tuning-based defense approaches. Notably, the defense method based on the most effective attack technique performs the best, reducing the attack success rate (ASR) to nearly zero in certain scenarios. Our contributions are summarized as follows:
@@ 4.2.4 Fake Completion Defense with Template
Table 1: The results of our defense methods compared to baselines against various attack methods in the direct prompt injection scenario. The evaluation metric used is ASR. Bold indicates the best performance. All results are reported in %.
@@ 5.1 Experimental Settings
We select popular and strong open-source LLMs as victim models for our experiments. Specifically, we choose Llama3.1-8b-Instruct Dubey et al. (2024), Qwen2-7b-Instruct Yang et al. (2024), and Llama3-8b-Instruct AI@Meta (2024). Throughout the experiments, unless otherwise specified, “Llama3”, “Llama3.1” and “Qwen2” refer to Llama3-8b-Instruct, Llama3.1-8b-Instruct, and Qwen2-7b-Instruct, respectively.
Evaluation Setups.
In our experimental setup, we assume that for our methods, only the utilized attack method is known during defense, and all other attack methods remain unknown. This setup challenges the generalization ability of our methods. For the security metric, we follow the evaluation protocol of Chen et al. (2024), using the attack success rate (ASR) to assess the effectiveness of the defense methods. We detecte if the answer to the injected instruction appears in the generated response. For the utility metric, we use accuracy to evaluate the potential negative impact of defense methods on model performance. Specifically, we employ the filtered QA dataset Li et al. (2023) and the sentiment analysis dataset SST2 Socher et al. (2013), which are not attacked and include the defense mechanism. We request the LLMs to answer the questions and verify whether the correct (golden) answers appear in the responses.
@@
Table 2: The results of our defense methods compared to baselines against various attack methods in the indirect prompt injection scenario. The evaluation metric used is ASR. Bold indicates the best performance. All results are reported in %.
@@ 5.3.1 Defense against Direct Attack
We perform the direct prompt injection attack following the approach of Chen et al. (2024), using 208 samples from AlpacaFarm. Table 1 presents the effectiveness of our defense methods in the direct prompt injection scenario. The results show that our methods, which are based on attack techniques, outperform the baselines, regardless of the attack method or the victim model. Among the baseline methods, the “Sandwich” method performs better on average than the others. The key difference between “Sandwich” and the other baselines lies in the position of the defense prompt: “Sandwich” places the defense prompt at the end of the data, similar to our methods. This suggests that placing the defense prompt at the end may interfere with the attack and enhance the defense’s effectiveness. When comparing the victim models, we find that Qwen2 is more vulnerable to the attacks, compared to the other two models.
@@ 5.3.2 Defense against Indirect Attack
In addition to evaluating defense against direct prompt injection attacks, we also assess its effectiveness against indirect attacks. The key difference between direct and indirect prompt injection attacks is that, in the case of indirect attacks, the input data is retrieved from external tools, such as search engines, and users are often unaware of the attack. To evaluate indirect prompt injection attacks, we use the filtered QA dataset from Li et al. (2023). Table 2 shows the results of our defense methods compared to the baselines in the indirect scenario. Our methods continue to outperform the baselines by a significant margin. When comparing both direct and indirect prompt injection attacks, it appears that indirect attacks are easier to defend against. Furthermore, Qwen2 remains the most susceptible model to attacks compared to the other two models.
@@ 5.3.3 Model Utility
A key evaluation metric for defense methods is their potential impact on the model’s utility. To assess the impact of our method, we use the filtered QA dataset from Li et al. (2023). For simplicity, we do not introduce any attacks into the retrieved data content, and we only verify whether the correct (golden) answer appears in the model’s response, with different defense methods. Table 3 presents the utility performance of various defense strategies. Notably, most defense strategies do not significantly affect the model’s utility. Moreover, our proposed defense methods can even improve the performance in some scenarios. Additionally, to further validate the robustness of our results, we conduct experiments on the sentiment analysis task using the SST2 dataset Socher et al. (2013), with results shown in Table 10. The results demonstrate that our methods cause minimal degradation to the model’s overall performance.
@@
Table 3: The general model performance on QA task, when applied with different defense methods. The evaluation metric is accuracy. The results are reported in %.
@@ 5.4 Ablation Study
In this section, we address several questions regarding our defense methods. We perform comprehensive experiments to solidify the validity and robustness of our approach.
@@
To further validate the effectiveness of our methods, we apply our methods to the closed-source models “GPT-3.5-Turbo” Jiang et al. (2023) and “GPT-4o-Latest” Hurst et al. (2024). Because we cannot change the conversation template, we only compare our methods based on “Ignore attack” and “Fakecom attack” with the baselines against direct prompt injection attack. Table 7 shows the results. From the table we can find out that our methods are also effective on closed-source models, surpassing the previous defense baselines. What’s more, comparing the defense performance of the two models with our defense methods reveals that stronger model is more suitable to our methods, making our methods more applicable.
@@
Beyond prompt-engineering-based attacks, we also evaluate the effectiveness of our defense methods against gradient-based attacks. Specifically, we perform direct prompt injection attacks using the GCG method Zou et al. (2023) and the AutoDAN method Zhu et al. (2023) with Llama3. Table 4 presents the defense results. Our first observation is that compared to baseline methods, our defense strategies more effectively mitigate these attacks. Notably, the “Fakecom-t” method proves to be the most effective, reducing the ASR to around 10% and demonstrating strong transferability across different attack types.
Defense Methods Attack-GCG Attack-AutoDAN
None 87.01 68.75
Sandwich 19.23 39.42
Instructional 28.84 52.88
Reminder 24.51 51.44
Isolation 40.38 54.32
Spotlight 19.71 24.51
Ours-Ignore 12.01 16.34
Ours-Escape 19.23 38.94
Ours-Fakecom 13.94 14.90
Ours-Fakecom-t 9.61 10.57
Table 4: The performance of the defense methods against the gradient-based attacks. The evaluation metric is ASR. Bold indicates the best performance. All results are reported in %.
@@
Although it’s very unlikely for the attacker to be aware of the conversation template, since application providers typically filter out template tokens, we are still interested in assessing the potential harm of such an attack. We utilize the direct prompt injection attack for evaluation and Table 5 presents the results. The table shows that the fake completion attack with a conversation template can be harmful, and most baseline methods are ineffective. Our methods, which rely on ignoring the attack and using the fake completion strategy, function as intended but result in only a limited decrease in ASR. Our method based on this attack (“Fakecom-t”) is effective, and this phenomenon raises our question: Would the effectiveness of the attack methods determine the effectiveness of defense methods designed on them?
@@
Table 5: The results show how harmful the fake completion attack with the conversation template is. The evaluation metric is ASR. The results are reported in %.
@@
In previous introduction, we argue that fine-tuning-based methods require significant computational resources. But the fine-tuning-based methods are more effective than previous prompt-engineering-based methods. Therefore, we compare our methods with StruQ Chen et al. (2024), which incorporates prompt injection attack methods into the clean data for fine-tuning. We incorporate the “Naive attack” and the “Ignore attack” respectively for evaluation. Table 6 shows the results. It’s obvious that the ability of StruQ to generalize to the unknown attacks is not satisfactory. Because “Ignore attack” is a part of “Combined attack”, “StruQ-Ignore” can defend against “Combined attack” successfully. The generalization ability of our methods is much better, effectively defending against different attacks.
Deal with long user input instructions.
When the user input instruction is long, our methods which append it at the end of the prompt may exceed the LLM’s context window. A potential solution is to truncate the original input instruction from the beginning of the prompt while retaining our defense prompt at the end. However, current benchmarks have not covered this problem. To assess the impact of this proposed approach, we conduct experiments with Llama3 in the direct scenario, by deleting original input instructions. The results, presented in Table 8, indicate that deleting the original input instruction has minimal impact on the defense performance of our methods. However, it significantly affects the baseline “Sandwich” method, highlighting the robustness of our defense methods. Additionally, we examine whether deleting the original input instruction affects the model’s general performance. Following the setup in Section 5.3.3, we conduct experiments with Llama3 on QA task. As shown in Table 9, this deletion does not degrade the model’s overall performance.
Defense Methods Naive Ignore Escape Fakecom Combined
None 10.55 53.35 88.25 75.30 86.00
StruQ-Naive 0.50 0.60 2.20 35.55 27.30
StruQ-Ignore 0.05 0.05 8.00 35.70 0.05
Ours-Ignore 0.05 0.35 0.30 0.10 1.35
Table 6: Defense performance of our methods and the fine-tuning method StruQ. “StruQ-Naive” means StruQ incorporates the “Naive attack” for fine-tuning. The evaluation metric is ASR. Bold indicates the best performance. All results are reported in %.
Impact of deleting data content.
We also examine the impact of deleting data content. For more details, please refer to Appendix A.3.
@@ 5.5 Case Study
Figure 16 provides two examples of responses without defense against the “Ignore attack” and “Fakecom attack”, both generated by Llama3. Additionally, we include the response generated by the model with defense method based on “Fakecom-t”, as it is the most effective defense approach. From the examples, we observe that the “Ignore attack” does not consistently persuade the LLM to ignore the original input instruction and the model may end up executing both instructions. Although in this instance, the “Fakecom attack” successfully misleads the LLM to execute the injected instruction directly, this strategy does not always work, and there are cases where the model executes both instructions, explaining the failures of the defense methods based on these attacks. In the case of the defense method based on “Fakecom-t”, we can observe that the defense method successfully enables the LLM to bypass the injected instruction. What’s more, the response still remains relevant to the original task, suggesting the defense method has little damage on the utility of the model.
@@ 6 Conclusion
Limitations
In this paper, we propose defense methods inspired by existing attack strategies. However, since a benchmark of long queries for prompt injection research has not yet been established, we are unable to conduct a thorough investigation into how the truncation method addresses the long-query problem, as discussed in ablation study. As an alternative, we remove the original input instructions from existing benchmarks and provide approximate results. These results demonstrate the effectiveness of the proposed methods. Moreover, we do not employ gradient-based attack methods as defense methods, as previous studies have shown that their performance is not satisfactory. Finally, since our methods are based on prompt engineering, we focus on conducting comprehensive experiments to demonstrate their effectiveness, rather than providing a mathematical proof to explain why they work. This limitation can also be found in other prompt injection studies Liu et al. (2024b); Chen et al. (2024); Li et al. (2023); Hines et al. (2024), regardless of whether they are fine-tuning-based or prompt-engineering-based.
Ethical Consideration
@@ Appendix A Appendix
A.1 Implementation Details.
A.2 Defense Baselines
@@
A.3 Impact of deleting data content
A straightforward approach to defending against indirect prompt injection attacks is to avoid retrieving data content altogether. To evaluate the impact of this strategy, we examine its effect on the QA task, assessing the LLMs’ ability to answer questions using only their inherent knowledge. We conduct experiments both with and without applying our defense methods, as shown in Table 11. The results indicate that retrieved data is essential, removing it significantly degrades the model’s performance.
A.4 Dataset Additional Information
@@
Defense Methods GPT-3.5-Turbo GPT-4o-Latest
None 32.69 50.48 32.69 88.46 87.50 65.86 92.78 63.46 100.00 100.00
Sandwich 13.94 17.30 8.65 4.32 42.30 15.86 29.32 10.09 5.76 37.98
Reminder 11.05 10.57 10.96 9.61 26.92 14.90 27.88 17.30 89.42 82.69
Isolation 22.59 39.42 24.51 43.26 77.40 52.40 83.17 52.88 94.71 99.51
Spotlight 16.34 31.73 13.46 15.38 71.15 19.71 45.67 15.38 47.11 68.75
Ours-Ignore 2.88 3.36 1.44 0.48 4.32 0.90 0.90 0.40 0.0 0.0
Ours-Fakecom 5.57 12.01 1.44 0.0 14.90 7.21 2.88 5.76 0.90 7.21
Table 7: The results of our defense methods compared with defense baselines applied on closed-source models. The evaluation metric is ASR. Bold indicates the best performance. All results are reported in %.
@@
None 89.90 87.02 93.75 89.90 75.00
Sandwich 44.23 65.38 34.61 41.34 61.53
Ours-Ignore 9.13 18.75 3.84 5.28 18.75
Ours-Escape 29.32 32.69 16.82 18.75 25.48
Ours-Fakecom 27.88 49.51 23.07 2.88 16.82
Ours-Fakecom-t 6.73 5.76 6.73 2.88 3.84
Table 8: Defense performance of our methods and baselines after deleting the original input instruction. The evaluation metric is ASR. Bold indicates the best performance. All results are reported in %.
Defense Methods QA Accuracy
None 78.05
Sandwich 79.90
Ours-Ignore 80.75
Ours-Escape 82.05
Ours-Fakecom 82.25
Ours-Fakecom-t 81.75
Table 9: QA accuracy on the Llama3 model when the original input instruction is deleted. The evaluation metric is accuracy. All results are reported in %. “None” refers to the standard input, where the original instruction remains unchanged, and no defense prompt is appended.
@@
None 94.83 94.15 94.06
Sandwich 95.29 94.49 95.64
Instructional 94.83 93.69 95.75
Reminder 94.72 93.34 96.67
Isolation 95.41 94.03 95.98
Spotlight 93.92 92.77 92.43
Ours-Ignore 95.41 95.41 93.46
Ours-Escape 95.98 93.23 92.66
Ours-Fakecom 95.18 94.61 92.54
Ours-Fakecom-t 95.98 94.83 95.29
Table 10: The general model performance on sentiment analysis task, when applied with different defense methods. The evaluation metric is accuracy. All the results are reported in %.
Defense Methods Llama3 Llama3.1 Qwen2
None 41.15 42.60 37.90
Ours-Ignore 42.10 42.00 38.55
Ours-Escape 40.45 39.50 37.45
Ours-Fakecom 40.80 40.50 37.95
Ours-Fakecom-t 43.40 43.40 37.10
Table 11: The results on QA task when the retrieved data is deleted. The evaluation metric is accuracy. All the results are reported in %.